OpenStudy (anonymous):

A horse pulls on an object with a force of 300 newtons and does 12,000 joules of work. How far was the object moved?

OpenStudy (anonymous):

@ShailKumar help please..?

OpenStudy (abhisar):

Hello @Miguel27525 !

OpenStudy (abhisar):

Lets do this

OpenStudy (abhisar):

WOrk = Force*distance

OpenStudy (abhisar):

Let the distance travelled be d meters, Now 12000=300*d -->d=12000/300=40m

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

Okayy thank you and im sorry i didnt reply all of the power went out and just came on

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)